A car radiator contains 10 liters of a 20 percent antifreeze solution. How many liters will have to be drained and then replaced with pure anitfreeze if the resulting solution is to be 50 percent antifreeze?
Answer by stanbon(75887)

A car radiator contains 10 liters of a 20 percent antifreeze solution. How many liters will have to be drained and then replaced with pure anitfreeze if the resulting solution is to be 50 percent antifreeze?
------
Equation:
anti - anti + anti = anti
0.20*10 - 0.20x + 1.00x = 0.50*10
-------
Multiply thru by 100 to get:
20*10 - 20x + 100x = 50*10
80x = 30*10
x = 3.75 liters (amt. that must be removed and replaced)
======================
Cheers,
Stan H.